The profit on a small appliance is given by P = x² -3x -60, in which x is the number of appliances sold per day. How many appliances were sold on a day when there was a $20 loss?
Answer by stanbon(48558)

The profit on a small appliance is given by P = x² -3x -60, in which x is the number of appliances sold per day. How many appliances were sold on a day when there was a $20 loss?
-----------------
A $20 loss means P = -20
x^2-3x-60 = -20
x^2-3x-40 = 0
(x-8)(x+5)=0
Positivesolution:
x = 8 (8 appliances were sold on the day $20 was lost)
=================
Cheers,
Stan H.